OpenStudy (anonymous):

solve for x: (x + 4)/3 = 2 a) x = -2 b) x = 2 c) x = 2/3 d) x = - 10/3

OpenStudy (anonymous):

Since this is multiple choice, try inserting the given numbers in for x. (a is out for sure)

OpenStudy (austinl):

\(\dfrac{x+4}{3}=2\) Multiply both sides by 3. \(x+4=6\) Subtract 4 from each side, \(x=6-4=~?\)

OpenStudy (anonymous):

so b?

OpenStudy (anonymous):

Lol @Niall XD Yes.

OpenStudy (anonymous):

CORRECT!

OpenStudy (anonymous):

ok, tank yew!

OpenStudy (anonymous):

Your Welcome! :)